OpenStudy (anonymous):

find th grapgh of (z+z_)^2+(z-z)^2=15+zz_

OpenStudy (chaise):

Are thee supposed to be complex numbers? With conjugates? Or what?

OpenStudy (anonymous):

Z=x+yi and z_=x-yi (imaginable numbers..!!

OpenStudy (chaise):

I thought so. Just substitute your values in for z and z_. (Imaginary numbers are complex number), and foil it out.

OpenStudy (anonymous):

i i have trie,and come up till..11x^2+19y^2=0 maybe en ellipse??